A child has near point at 10 cm. What is the maximum angular magnification the child can have with a convex lens of focal length 10 cm?

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. The child has D = 10 cm and f = 10 cm The maximum angular magnification is obtained when the image is formed at near point. m = 1 + D/f = 1 + 10/10 = 1 + 1 = 2